A group of students got the following scores in an achievement test:9,12,15,18,21 and 24 . consider the sample of size 3 that can be drawn from this population. Construct a sampling distribution of the resulting means and find the probability




1
Expert's answer
2022-06-07T09:29:22-0400

We have population values 9,12,15,18,21, 24, population size N=6 and sample size n=3.

Mean of population "(\mu)" = "\dfrac{9+12+15+18+21+24}{6}=16.5"

Variance of population 



"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n}=\dfrac{1}{6}(56.25+20.25+2.25""+2.25+20.25+56.25)=26.25""\sigma=\sqrt{\sigma^2}=\sqrt{26.25}\approx5.1235"

Select a random sample of size 3 without replacement. We have a sample distribution of sample mean.

The number of possible samples which can be drawn without replacement is "^{N}C_n=^{6}C_3=20."

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 & 9,12,15 & 12 \\
 \hdashline
 2 & 9,12,18 & 13 \\
 \hdashline
 3 & 9,12,21 & 14 \\
 \hdashline
 4 & 9,12,24 & 15 \\
 \hdashline
 5 & 9,15,18 & 14 \\
 \hdashline
 6 & 9,15,21 & 15 \\
 \hdashline
 7 & 9,15, 24 & 16 \\
 \hdashline
 8 & 9,18,21 & 16 \\
 \hdashline
 9 & 9,18,24 & 17 \\
 \hdashline
 10 & 9, 21,24 & 18 \\
 \hdashline
 11 & 12,15,18 & 15 \\
 \hdashline
 12 & 12, 15,21 & 16 \\
 \hdashline
 13 & 12, 15, 24 & 17 \\
 \hdashline
 14 & 12,18,21 & 17 \\
 \hdashline
 15 & 12,18,24 & 18 \\
 \hdashline
 16 & 12, 21,24 & 19 \\
 \hdashline
 17 & 15, 18,21 & 18 \\
 \hdashline
 18 & 15, 18,24 & 19 \\
 \hdashline
 19 & 15, 21,24 & 20 \\
 \hdashline
 20 & 18, 21,24 & 21 \\
 \hdashline
\end{array}"






"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X}) & \bar{X}^2f(\bar{X})
\\ \hline
 12 & 1/20 & 12/20 & 144/20 \\
 \hdashline
 13 & 1/20 & 13/20 & 169/20 \\
 \hdashline
 14 & 2/20 & 28/20 & 392/20 \\
 \hdashline
 15 & 3/20 & 45/20 & 675/20 \\
 \hdashline
 16 & 3/20 & 48/20 & 768/20 \\
 \hdashline
 17 & 3/20 & 51/20 & 867/20 \\
 \hdashline
 18 & 3/20 & 54/20 & 972/20 \\
 \hdashline
 19 & 2/20 & 38/20 & 722/20 \\
 \hdashline
 20 & 1/20 & 20/20 & 400/20 \\
 \hdashline
 21 & 1/20 & 21/20 & 441/20 \\
 \hdashline
\end{array}"



Mean of sampling distribution 


"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=\dfrac{330}{20}=16.5=\mu"



The variance of sampling distribution 


"Var(\bar{X})=\sigma^2_{\bar{X}}=\sum\bar{X}_i^2f(\bar{X}_i)-\big[\sum\bar{X}_if(\bar{X}_i)\big]^2""=\dfrac{5550}{20}-(16.5)^2=5.25= \dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})""\sigma_{\bar{X}}=\sqrt{5.25}\approx2.2913"